OpenStudy (anonymous):

Daniel is using logarithms to solve the equation 24x = 14. Which of the following equations would be equivalent to his original expression?

OpenStudy (anonymous):

These are the choices 2 log 4x = log 14 4x log 2 = log 14 2 log 4 = x log 14 x log 4 = 2 log 14

jimthompson5910 (jim_thompson5910):

The original equation is \[\Large 2^{4x} = 14\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@Isaih.Feynman?

jimthompson5910 (jim_thompson5910):

Take the log of both sides. Then use rule 3 on this page http://www.purplemath.com/modules/logrules.htm to get \[\Large \log(2^{4x}) = \log(14)\] \[\Large 4x\log(2) = \log(14)\]

OpenStudy (anonymous):

Thank you for helping jim

OpenStudy (anonymous):

Thank you so very much Jim! I was so desperate for that answer. I am taking a 50 question test and that was the only one I was missing! I just made this account and I don't know how to give medals or anything, but thank you!